### v1

@@ 9 Averaging observations
For weights, we’ll use inverse square weighting in a measure of the interval size. The individual confidence intervals may not be symmetric about the MLEs, and there are several ways we could address this. For the sake of the example, we’ll use the (conservative) procedure of symmetrizing the interval, taking the maximum deviation from the MLE as our error bar. That is, we let
@@
It remains to be determined whether this in fact yields correct covereage.
To study the coverage of our averaged result, we simulate a simple situation. As above, we suppose we make several, $K$, observations of $\hat{\Gamma}_{k}=n_{k}/T_{k},k=1,\ldots,K$ with corresponding 68% confidence intervals $(\ell_{k},u_{k})$. We apply our weighted averaging procedure and ask how often the interval $\bar{\Gamma}\pm\bar{\delta}$ includes $\Gamma$. It will be sufficient for our demonstration to choose $\Gamma=1$ and $T_{k}=2,\forall k$. We’ll use the Garwood 68% confidence intervals for $(\ell_{k},u_{k})$ and consider averages of up to 50 observations. Figure 18 shows the result for coverage as a function of the number of observations being averaged. We show the target (0.68) coverage as a horizontal line; values above this correspond to overcovering and values below to undercovering.
@@
Even though the single measurement confidence intervals overcover (coverage probability 0.85 for our symmetrized intervals; note that this is quite a bit higher than the coverage of 0.72 for the Garwood interval for $\mu=2$) for this example, the coverage decreases as more results are averaged, and eventually intervals that severely undercover are obtained. In spite of the conservatism of our Poisson interval, the averaging process fails to maintain frequency coverage. This may seem counter to expectations based on the central limit theorem. However, we can’t expect the central limit theorem to apply simply because we make a large number of observations. We also need to know the variances, which we don’t. It may be noted that doing the same simulation with normal samplings with known variance gives the expected 68% coverage for all the averages.
@@ 10 Discussion
We have investigated several of the most commonly used methods for quoting confidence intervals for the mean of a Poisson distribution, possibly including a background contribution. Even demanding “exactness” there is still a large amount of choice in algorithms. We wish to reach some conclusion among the trade-offs and arrive at a recommendation. We embark on a trip through our “desirable properties” of Section 3 to sort this out.
@@ 10.2 Connectedness
We have already argued that for the Poisson distribution, connected intervals are the sensible choice. Since if we start with a disconnected interval we can always fill in the values in the gap, all of our methods allow for this property and satisfy it as implemented here.
@@ 10.3 Contains maximum likelihood estimator
We argued (as does [21]) that an acceptable descriptive interval should contain the point estimator, i.e., the maximum likelihood estimator. The Crow&Gardner intervals are not guaranteed to include the MLE, e.g., Fig. 8b. We regard this as a serious flaw, although it is mitigated somewhat by appearing only at low confidence levels. The FC intervals also may not include the MLE. In this case, it is a design feature that $\theta<0$ is excluded and the MLE is redefined to never be quoted below zero. We argue that such a redefinition does not well serve the goal of description. While the frequency interpretation is preserved, this does limit its usefulness as a descriptive statistic.
@@ 10.4 Optimal coverage
The original Garwood interval has been much criticized for substantial overcoverage. Thus, a large amount of the statistics literature has been directed at inventing algorithms that minimize the coverage without undercovering. We have investigated several of the resulting algorithms. All of them have regions of parameter space with significant overcoveage, while doing better than the Garwood intervals.
From Fig. 4 (for 68% confidence level) we find that the Crow&Gardner intervals have substantially better coverage than the Garwood intervals. It also has better covereage in general than the Sterne intervals (Fig. 3) and the likelihood ratio intervals (Fig. fig:LRCrow). As the score and likelihood ratio intervals have similar coverage (Fig. fig:scoreLR), the Crow&Gardner intervals also improve coverage over the score intervals. The comparison of coverage for Crow&Gardner and Kabaila&Byrne intervals (Fig. 9) is somewhat more mixed, with one or the other performing better depending on $\mu$. However, the general impression favors Crow&Gardner once again. The Blaker intervals perform somewhat better than the Sterne intervals (Fig. fig:SternBlaker), but not so much as to make them better than Crow&Gardner. Thus, Crow&Gardner intervals appear to be among the best in terms of covereage. We note that the Garwood interval only satisfies Eq. 19 asymptotically (besides perhaps some fortuitous solutions).
@@ 10.5 Length
Traditionally (e.g., [24, 20, 21, 40, 25]), statisticians have focused on optimizing length as well as covereage. To compare algorithms for length, we consider the expected length of an interval, as a function of $\theta+b$. For our conventional methods, this is independent of $b$, so we may simply consider the expected length as a function of $\theta$ for $b=0$. For any given algorithm, giving confidence interval $(\theta_{\rm min},\theta_{\rm max})$, this length is
@@
This is shown in Fig. 19 for several algorithms.
Figure 19: Left: Expected length of 68% CL intervals, as a function of $\theta$ (assuming $b=0$). Right: Same as left figure, except divided by $2\sqrt{\theta}$, for $\theta$ between 0.1 and 5.
The detailed behavior of the lengths in Fig. 19 depend slightly on the value of $\theta$. For $\theta\sim 1$, The ordering from largest to smallest lengths is: Garwood, Sterne, Blaker, score, likelihood ratio, and Crow&Gardner. The Kabaila&Byrne intervals are also somewhat larger than the Crow&Gardner intervals (Fig. 9).
@@
On the two traditional measures of performance, coverage and length, the Crow&Gardner intervals look very attractive, while the original Garwood intervals come in last. However, among nested equal-tailed exact intervals, Thulin and Zwanzig [18] show that the Garwood interval minimizes the expected interval length for all $\theta$ and minimizes the length for all $N=n$.
@@ 10.6 Scale and asymptopia
It is probably not controversial that in the limit of large statistics, the confidence intervals should converge on the correct coverage. Fortunately, (e.g., see the right plot in Fig. 19), all of our exact (and $\sqrt{n}$ and Bayes) intervals satisfy this requirement.
More subtle is our notion that a good descriptive interval should have an intuitive scale. We know that the variance of a Poisson with mean $\mu$ is $\mu$, hence our intervals should scale approximately as $\sqrt{N}$ as remarked in Section 3.6. From Figs. 2-4, 6, and 9-11, we find that most of our intervals satisfy this intuition. The two Bayes intervals (Fig. 13) also satisfy it. The notable exceptions are the FC intervals and the cut likelihood ratio interval in the presence of background, (Fig. 16 and Fig. 17). In these cases, the confidence intervals are much smaller than the desired intuitive behavior when negative fluctuations occur at low statistics. That is, these intervals suggest a precision in the measurement much better than actual.
@@ 10.7 Ordered
In fact, not all methods have monotonically increasing interval sizes with observation $n$. The Sterne and Blaker intervals violate this, as can be seen in the plot on the right in Fig. 6, even though they satisfy strict ordering. Intervals derived from the likelihood ratio test (hence also FC) also have intervals that sometimes decrease in size as $n$ increases. Score intervals are non-decreasing but sometimes have the same length for successive $n$ values. From Fig. 9, we see that the Kabaila&Byrne intervals rather badly violate monoticity of interval size as a function of observation. This is interesting, because Kabaila&Byrne make a point of demanding strict ordering. It seems that in satisfying strict ordering this may have affected the monoticity of the interval length. We suggest that monoticity of interval length is more important than strict ordering, at least if ordering is satisfied.
@@ 10.8 Symmetry
The Garwood intervals are defined to be symmetric (in probability), enforcing $1-\alpha/2$ probability on each side. We have compared the Crow&Gardner symmetry in Fig. 5 and the Blaker in Fig. 7. In both cases, the excursions from symmetry are smaller for the Garwood intervals, as expected.
@@ 10.9 Nested
As defined in Section 3.9, a nested interval is one in which the interval for a smaller confidence level is always a subset of the interval for a larger confidence level. Strictly nested intervals are when this is always a proper subset. We have already seen some signs of issues with nestedness in Fig. 8b which shows that the Blaker intervals are nested, but not strictly nested, and the Crow&Gardner intervals are not even nested. In contrast, the Garwood interval is strictly nested, e.g., Fig. 21.
To investigate the behavior of the non-Garwood intervals, Thulin and Zwanzig [18] define the notion of a “strictly two-sided” test and corresponding strictly two-sided confidence intervals. The key point in this definition is that computing a $p$ value based on the test statistic requires simultaneously comparing the test statistic to both tails under the null distribution. The Garwood interval is not such a test, because the $p$ value for $\theta=\theta_{0}$ is
@@
In this case, the $p$ value is computed as twice the minimum of the two tails. This is consistent with our intuitive notion that a $p$ value is the probability (under $H_{0}$) of obtaining a sampling as far or farther from the null hypothesis than the observed value. On the other hand, Thulin and Zwanzig show that the likelihood ratio, Sterne, score, and Blaker intervals are all strictly two-sided intervals. They refer to the Crow&Gardener and Kabaila&Byrne intervals as “strictly two-sided type”, because they are not derived from test statistics, but have similarities with the strictly two-sided intervals. Strictly two-sided tests are found to have the following properties [18]:
@@
    The $p$ value is a discontinuous function of the null hypothesis ($\theta_{0}$);
@@
Hence, the lack of strict nesting or even nesting behavior we see for example in Fig. 8b. The likelihood ratio, Sterne, and score intervals also are not strictly nested and the Kabailla&Byrne intervals are not nested.
@@
Figure 21: The Garwood confidence interval as a function of confidence level for $n=2$.
10.10  Continuity
@@
Continuity (Section 3.10) here refers to the continuity of the confidence intervals as a function of confidence level. Even if the intervals are nested, if they are not continuous a small change in confidence level could correspond to a large change in interval. The Garwood intervals are continuous (e.g., Fig. 21).
@@
Similarly with nestedness, Thulin&Zwanzig [18] find that strictly two-sided intervals are discontinuous functions of confidence level. We see an example of this in Fig. 8b for the Blaker interval. We also see that the strictly two-sided type interval, the Crow&Gardner is not even a monotonic function of confidence level. This is also true for the Kabailla&Byrne strictly two-sided type interval. In this case, increasing the confidence level can actually decrease the interval.
10.11  Gives sensible $p$ values
@@
According to the theorem quoted in Section 10.9, $p$ values based on the strictly two-sided tests are discontinuous functions of the null hypothesis $\theta=\theta_{0}$. For example, the left side of Figure 22 shows a comparison of the $p$ values derived from the Garwood and the likelihood ratio confidence intervals, as a function of null hypothesis $\theta_{0}$, for an observation of $n=3$.
@@
Figure 22: The $p$ values as a function of null hypothesis $\theta_{0}$ for an observation $n=3$ (and no background). Left: The continuous orange curve is according to the Garwood interval and the discontinuous purple curve is according to the likelihood ratio interval. Right: The $p$ values according to the Kabailla&Byrne intervals, also for $n=3$.
@@
Overall, the Garwood $p$ values tend to be larger, indicating less discrimination against the null hypothesis. This is a reflection of the fact that the Garwood intervals tend to overcover more, which is the major criticism of these intervals.
@@
On the other hand, the Garwood $p$ values are well-behaved, being continuous and monotonically rising to a plateau of $p=1$ when $\theta_{0}$ is near $n=3$, and then monotonically falling when $\theta_{0}$ is somewhat larger than three. The Garwood $p$ values are “bimonotonic decreasing”, in the above sense. In contrast, the behavior of the likelihood ratio $p$ values is much more problematic. First, they are discontinuous, so that a small change in null hypothesis ($\theta_{0}$) can give a large change in $p$, which is not something we would normally expect given the continuity of our probability distribution as a function of $\theta$. Second, there is a region where the $p$ value becomes higher for a more inconsistent value of $\theta_{0}$, contrary to anticipated reasonable behavior. The strictly two-sided type intervals are no better. For example, the right side of Fig. 22 shows the $p$-values for $n=3$ computed according to the Kabailla-Byrne intervals. Now the $p$ values are not even well-defined, taking on multiple values for many values of $\theta_{0}$.
@@
We regard these issues as serious flaws in a $p$ value. A discontinuous or non-bimonotonic $p$ value is not expected behavior, and provides a non-intuitive description of the measurement. Hirji [16] states that continuous bimonotonic decreasing behavior in the $p$ values implies some desirable properties: (i) A two-sided test which is significant at $\alpha_{1}$ is also significant at $\alpha_{2}$ if $\alpha_{2}>\alpha_{1}$; (ii) The confidence intervals are nested; (iii) The confidence intervals are connected. Thulin and Zwanzig [18] show further that if the limits of a confidence interval are continuous in $\alpha$ then the derived $p$ value function is strictly bimonotone in $\theta$.
@@
Only the Garwood intervals provide for sensible $p$ values among the two-sided exact intervals studied. Further, as noted earlier, the Garwood interval is optimal among nested equal-tailed exact intervals in the sense that it minimizes the expected interval length for all $\theta$ and minimizes the length for all $N=n$ [18] .

### v2

@@ 10 Averaging observations
For weights, we will use inverse square weighting in a measure of the interval size. The individual confidence intervals may not be symmetric about the MLEs, and there are several ways we could address this. For the sake of the example, we will use the (conservative) procedure of symmetrizing the interval, taking the maximum deviation from the MLE as our error bar. That is, we let
@@
It remains to be determined whether this in fact yields correct coverage.
To study the coverage of our averaged result, we simulate a simple situation. As above, we suppose we make several, $K$, observations of $\hat{\Gamma}_{k}=n_{k}/T_{k},k=1,\ldots,K$ with corresponding 68% confidence intervals $(\ell_{k},u_{k})$. We apply our weighted averaging procedure and ask how often the interval $\bar{\Gamma}\pm\bar{\delta}$ includes $\Gamma$. It will be sufficient for our demonstration to choose $\Gamma=1$ and $T_{k}=2,\forall k$. We will use the Garwood 68% confidence intervals for $(\ell_{k},u_{k})$ and consider averages of up to 50 observations. Figure 19 shows the result for coverage as a function of the number of observations being averaged. We show the target (0.68) coverage as a horizontal line; values above this correspond to overcovering and values below to undercovering.
@@
Even though the single measurement confidence intervals overcover (coverage probability 0.85 for our symmetrized intervals; note that this is quite a bit higher than the coverage of 0.72 for the Garwood interval for $\mu=2$) for this example, the coverage decreases as more results are averaged, and eventually intervals that severely undercover are obtained. In spite of the conservatism of our Poisson interval, the averaging process fails to maintain frequency coverage. This may seem counter to expectations based on the central limit theorem. However, we cannot expect the central limit theorem to apply simply because we make a large number of observations. We also need to know the variances, which we do not. It may be noted that doing the same simulation with normal samplings with known variance gives the expected 68% coverage for all the averages.
@@ 11 Discussion
We have investigated several of the most commonly used methods for quoting confidence intervals for the mean of a Poisson distribution, possibly including a background contribution. Even demanding “exactness” there is still a large amount of choice in algorithms. We wish to reach some conclusion among the trade-offs and arrive at a recommendation. We embark on a trip through our “desirable properties” of section 4 to sort this out, noting again that they may not all be achieved simultaneously.
@@ 11.2 Connectedness
We have already argued that for the Poisson distribution, connected intervals are the sensible choice. If we start with a disconnected interval we can always fill in the values in the gap(s). All of our methods allow for this property and satisfy it as implemented here. We can improve coverage by relaxing this requirement, but this would be at the cost of unintuitive complexity and we do not consider doing this.
@@ 11.3 Contains maximum likelihood estimator
We argued (section 4.3, as does [6]) that an acceptable descriptive interval should contain the point estimator, i.e., the maximum likelihood estimator. The Crow&Gardner intervals are not guaranteed to include the MLE, e.g., Fig. 8b. While we generally regard this as a serious flaw, in this case it appears only at low confidence levels, below those usually used in quoting confidence intervals. Hence, it may not be an issue in practice. The FC intervals also may not include the MLE, in our usage of the term. However, it is a design feature of FC intervals that $\theta<0$ is excluded and the MLE is defined to never be quoted below zero [8]. Hence, in this sense, the FC intervals do include the MLE. We have argued in section 3 that such a restriction does not well serve the goal of description.
@@ 11.4 Optimal coverage
The original Garwood interval has been much criticized for substantial overcoverage. Thus, a large amount of the statistics literature has been directed at inventing algorithms that minimize the coverage without undercovering. We have investigated several of the resulting algorithms. All of them have regions of parameter space with significant overcoverage, while doing better than the Garwood intervals.
From Fig. 4 (for 68% confidence level) we find that the Crow&Gardner intervals have substantially better coverage than the Garwood intervals. It also has better coverage in general than the Sterne intervals (Fig. 3) and the likelihood ratio intervals (Fig. 10). As the score and likelihood ratio intervals have similar coverage (Fig. 11), the Crow&Gardner intervals also improve coverage over the score intervals. The comparison of coverage for Crow&Gardner and Kabaila&Byrne intervals (Fig. 9) is somewhat more mixed, with one or the other performing better depending on $\mu$. However, the general impression favors Crow&Gardner once again. The Blaker intervals perform somewhat better than the Sterne intervals (Fig. 6), but not so much as to make them better than Crow&Gardner. Thus, Crow&Gardner intervals appear to be among the best in terms of coverage. We note that the Garwood interval only satisfies Eq. 20 asymptotically (besides perhaps some fortuitous solutions).
However, as will be discussed in sections 11.711.11 and in the Conclusions, section 12, we must consider the incompatibility of the methods with smallest coverage with other desirable properties, and will decide not to insist on optimal coverage.
@@ 11.5 Length
Traditionally (e.g., [41, 10, 6, 51, 40]), statisticians have focused on optimizing length as well as coverage. To compare algorithms for length, we consider the expected length of an interval, as a function of $\theta+b$. For our conventional methods, this is independent of $b$, so we may simply consider the expected length as a function of $\theta$ for $b=0$. While we discuss other measures of length in section 4.5, the expected length as a function of $\theta$ is sufficient for our discussion here. For any given algorithm, giving confidence interval $(\ell(n),u(n))$, this length is
@@
This is shown in Figs. 20 and 21 for several algorithms.
Figure 20: Expected length of 68% CL intervals, as a function of $\theta$ (assuming $b=0$)..
Figure 21: Same as Fig. 20, except divided by $2\sqrt{\theta}$, and for $\theta$ between 0.1 and 5.
The detailed behavior of the lengths in Figs. 20 and 21 depend slightly on the value of $\theta$. For $\theta\sim 1$, The ordering from largest to smallest lengths is: Garwood, Sterne, Blaker, score, likelihood ratio, and Crow&Gardner. The Kabaila&Byrne intervals are also somewhat larger than the Crow&Gardner intervals (Fig. 9).
@@
On the two traditional measures of performance, coverage and length, the Crow&Gardner intervals look very attractive, [^16] At one point I thought that the Crow&Gardner intervals would be the recommended choice, before I realized the substantial issues with other considerations. while the original Garwood intervals come in last. However, among strictly nested equal-tailed exact intervals, Thulin and Zwanzig [1] show that the Garwood interval minimizes the expected interval length for all $\theta$ and minimizes the length for all $N=n$.
@@ 11.6 Scale and asymptopia
It is probably not controversial that in the limit of large statistics (large $\mu$), the confidence intervals should converge on the correct coverage. Fortunately, (e.g., see Fig. 21), all of our exact (and $\sqrt{n}$ and Bayes) intervals satisfy this requirement, though the convergence of the Garwood seems slowest, reflecting its overall length performance.
More subtle is our notion that a good descriptive interval should have an intuitive scale, reflecting the measurement precision. We know that the variance of a Poisson with mean $\mu$ is $\mu$, hence our intervals should scale approximately as $\sqrt{N}$ as remarked in section 4.6. From Figs. 2-4, 6, and 9-11, we find that most of our intervals satisfy this intuition. The two Bayes intervals (Fig. 13) also satisfy it. The notable exceptions are the FC intervals and the cut likelihood ratio interval in the presence of background, (Fig. 16 and Fig. 17). In these cases, the confidence intervals are much smaller than the desired intuitive behavior when negative fluctuations occur at low statistics. That is, these intervals suggest a precision in the measurement much better than actual. FC [13] suggest quoting the “sensitivity” of the experiment as a way to alert the consumer when this occurs.
@@ 11.7 Ordered
In fact, not all methods have monotonically increasing interval sizes with observation $n$. The Sterne and Blaker intervals violate this, as can be seen in the plot on the right in Fig. 6, even though they satisfy strict ordering. Intervals derived from the likelihood ratio test (hence also FC) also have intervals that sometimes decrease in size as $n$ increases. Score intervals are non-decreasing but sometimes have the same length for successive $n$ values. From Fig. 9, we see that the Kabaila&Byrne intervals rather badly violate monotonicity of interval size as a function of observation. This is interesting, because Kabaila&Byrne make a point of demanding strict ordering. It seems that in satisfying strict ordering this may have affected the monotonicity of the interval length. We suggest that monotonicity of interval length is more important than strict ordering, at least if ordering is satisfied.
@@ 11.8 Symmetry
The Garwood intervals are defined to be symmetric (in probability), enforcing $1-\alpha/2$ probability on each side. We have compared the Crow&Gardner symmetry in Fig. 5 and the Blaker in Fig. 7. In both cases, the excursions from symmetry are smaller for the Garwood intervals, as expected. We do not necessarily consider optimizing on symmetry to be crucial, but it is intuitively helpful to have a description where we know that the probability is reasonably balanced between the two tails.
@@ 11.9 Nested
As defined in Section 4.9, a nested interval is one in which the interval for a smaller confidence level is always a subset of the interval for a larger confidence level. Strictly nested intervals are when this is always a proper subset. We have already seen some signs of issues with nestedness in Fig. 8 which shows that the Blaker intervals are nested, but not strictly nested, and the Crow&Gardner intervals are not even nested. In contrast, the Garwood interval is strictly nested, e.g., Fig. 23.
To investigate the behavior of the non-Garwood intervals, Thulin and Zwanzig [1] define the notion of a “strictly two-sided” test and corresponding strictly two-sided confidence intervals. The key point in this definition is that computing a $p$-value based on the test statistic requires simultaneously comparing the test statistic to both tails under the null distribution. The Garwood interval is not such a test, because the $p$-value for $\theta=\theta_{0}$ is
@@
In this case, the $p$-value is computed as twice the minimum of the two tails. This is consistent with our intuitive notion that a $p$-value is the probability (under $H_{0}$) of obtaining a sampling as far or farther from the null hypothesis than the observed value. On the other hand, Thulin and Zwanzig show that the likelihood ratio, Sterne, score, and Blaker intervals are all strictly two-sided intervals. They refer to the Crow&Gardner and Kabaila&Byrne intervals as “strictly two-sided type”, because they are not derived from test statistics, but have similarities with the strictly two-sided intervals. Strictly two-sided tests are found to have the following properties [1]:
@@
    The $p$-value is a discontinuous function of the null hypothesis ($\theta_{0}$);
@@
Hence, the lack of strict nesting or even nesting behavior we see for example in Fig. 8. The likelihood ratio, Sterne, and score intervals also are not strictly nested and the Kabaila&Byrne intervals are not nested.
@@
Figure 23: The Garwood confidence interval $(\ell,u)$ as a function of confidence level for $n=2$. The upper curve is the upper bound $u$ and the lower curve is the lower bound $\ell$.
11.10 Continuity and monotonicity
@@
Continuity (Section 4.10) here refers to the continuity of the confidence intervals as a function of confidence level. We will discuss $p$-values in section 11.11. Even if the intervals are nested, if they are not continuous a small change in confidence level could correspond to a large change in interval. The Garwood intervals are continuous (e.g., Fig. 23).
@@
Similarly with the observations about nestedness, Thulin&Zwanzig [1] find that strictly two-sided intervals are discontinuous functions of confidence level. We see an example of this in Fig. 8(left) for the Blaker interval. It seems that still worse than discontinuity is lack of monoticity. We see that the strictly two-sided type interval, the Crow&Gardner is not even a monotonic function of confidence level (Fig. 8(right)). This is also true for the Kabaila&Byrne strictly two-sided type interval. In such cases, increasing the confidence level can actually decrease the interval, giving seriously counterintuitive behavior.
We note that some intervals do not have lengths that are monotonically increasing with $n$ (e.g., the Kabaila&Byrne intervals are especially notable in Fig. 9 and less so for the LR and score intervals in Fig. 11 and for the FC intervals in Fig. 18). While this is not desirable, we consider this to be less important than some of the other considerations.
11.11 Gives sensible $p$-values
@@
According to the theorem quoted in Section 11.9, $p$-values based on the strictly two-sided tests are discontinuous functions of the null hypothesis $\theta=\theta_{0}$. For example, the left side of Figure 24 shows a comparison of the $p$-values derived from the Garwood and the likelihood ratio confidence intervals, as a function of null hypothesis $\theta_{0}$, for an observation of $n=3$.
@@
Figure 24: The $p$-values as a function of null hypothesis $\theta_{0}$ for an observation $n=3$ (and no background). Left: The continuous orange curve is according to the Garwood interval and the discontinuous purple curve is according to the likelihood ratio interval. Right: The $p$-values according to the Kabaila&Byrne intervals, also for $n=3$.
@@
Overall, the Garwood $p$-values tend to be larger, indicating less discrimination against the null hypothesis. This is a reflection of the fact that the Garwood intervals tend to overcover more, which is the major criticism of these intervals.
@@
On the other hand, the Garwood $p$-values are well-behaved, being continuous and monotonically rising to a plateau of $p=1$ when $\theta_{0}$ is near $n=3$, and then monotonically falling when $\theta_{0}$ is somewhat larger than three. The Garwood $p$-values are “bimonotonic decreasing”, in the above sense. In contrast, the behavior of the likelihood ratio $p$-values is much more problematic. First, they are discontinuous, so that a small change in null hypothesis ($\theta_{0}$) can give a large change in $p$, which is not something we would normally expect given the continuity of our probability distribution as a function of $\theta$. Second, there is a region where the $p$-value becomes higher for a more inconsistent value of $\theta_{0}$, contrary to anticipated reasonable behavior. The strictly two-sided type intervals are no better. For example, the right side of Fig. 24 shows the $p$-values for $n=3$ computed according to the Kabaila-Byrne intervals. Now the $p$-values are not even well-defined, taking on multiple values for many values of $\theta_{0}$.
@@
We regard these issues as serious flaws in a $p$-value. A discontinuous or non-bimonotonic $p$-value is not expected behavior, and provides a non-intuitive description of the measurement. Hirji [17] states that continuous bimonotonic decreasing behavior in the $p$-values implies some desirable properties: (i) A two-sided test which is significant at $\alpha_{1}$ is also significant at $\alpha_{2}$ if $\alpha_{2}>\alpha_{1}$; (ii) The confidence intervals are nested; (iii) The confidence intervals are connected. Thulin and Zwanzig [1] show further that if the limits of a confidence interval are continuous in $\alpha$ then the derived $p$-value function is strictly bimonotone in $\theta$.
@@
Only the Garwood intervals provide for sensible $p$-values among the two-sided exact intervals studied. Further, as noted earlier, the Garwood interval is optimal among nested equal-tailed exact intervals in the sense that it minimizes the expected interval length for all $\theta$ and minimizes the length for all $N=n$ [1] . At least for the tests studied, the lack of strict nesting comes along with intervals that are not strictly momotonic in $\alpha$ and $p$-values that are discontinuous in the null hypothesis.